The mscrep Class for L^{A}TEX 2ε

Martin Reddy

April 25, 1996

1 Introduction

This document describes the mscrep class which provides a document style for presenting M.Sc. project reports in the department of Computer Science. To use this class, all you need do is begin your L^{A}TEX document with the declaration:

     \documentclass[<options>]{mscrep}

The default behaviour of this formatting class is to produce documents which are typeset as 10 point doubled-spaced text on single-sided A4 pages.

The reader’s attention is also drawn to the csthesis style file which provides a generic document style for submitting theses in the Computer Science department which are in line with the official University regulations (see “texdoc csthesis” for further details).

2 Class Options

Class options are specified in the the square-bracketed section of the \documentclass command. The mscrep class provides one class option, and allows only a couple of the standard L^{A}TEX 2ε report class options to be passed (e.g. leqno and fleqn). The following class option is currently supported by mscrep:

parskip : alters the paragraph formatting so that each paragraph is separated by a vertical space, and there is no indentation at the start of each paragraph. (The list environments are also updated to reflect this change.)

3 The Title Page

The mscrep class produces a title page which is formatted so that it may be bound using the standard departmental (orange) report covers. When this is the case, the project’s title, author and submission date will be visible, along with the header “M.Sc. Project Report”.

The following short example illustrates how to produce a title page using the mscrep class. Note that the abstract is displayed at the bottom of the title page and so must be defined before executing a \maketitle command.

  \title{<report_title>}
  \author{<student's_name>}
  \date{<date_of_submission>}
  \abstract{<report_abstract>}
  \maketitle

Each of the above four fields can include a \thanks{<text>} attribution to provide any additional comments (these will be formatted as a footnote at the bottom of the title page).